# Training Large Neural Networks With Low-Dimensional Error Feedback

### v3

@@ 3.2 Training the feedback weights
From the previous analysis, it is apparent that the feedback matrix $B$ must be trained for the learning dynamics to converge to the correct solution of Eq. (6). When $B$ is low-rank and fixed, the network lacks sufficient capacity to transmit the full error information, which can impede learning.
One approach is to adopt a learning rule inspired by Kolen-Pollack, as in eq. (2). When $B=QP$, the updates are given by
$\Delta W_{2}^{\mu}=\eta\boldsymbol{\delta}^{\mu}\boldsymbol{x}^{\mu T}W_{1}^{T}-\lambda W_{2}\quad\text{ and }\quad\Delta Q^{\mu}=\eta W_{1}\boldsymbol{x}^{\mu}\boldsymbol{\delta}^{\mu T}P^{T}-\lambda Q,$ (7)
where $\boldsymbol{\delta}^{\mu}=\frac{\partial\mathcal{L}(\boldsymbol{x}^{\mu},\boldsymbol{y}^{\mu})}{\partial\boldsymbol{y}}=\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu}$ is the error gradient for the $\mu$-th data point, $\eta$ is the learning rate, and $\lambda$ is the regularization parameter. Importantly, in this framework only the column space of $B$, or in our case $Q$ are updated.
However, adopting this learning framework is insufficient when $P$ remains fixed and randomly initialized (Figure 2c). Since $P$ can project onto at most $r$ unique directions of the error, it may not align with the relevant error subspace. In this case, the network may converge to an incorrect solution, as indicated by the non-uniqueness of solutions to (6) when $m>r$.
Ideally, we want $P$ to be an orthogonal matrix whose $r$ columns span the top $r$ principal directions of the output-output correlation matrix $\Sigma^{yy}$. This alignment ensures that the most significant components of the error are propagated back through the network.
In cases where the output correlations are unknown, we can update $P$ using a modified Oja learning rule [18]:
$\Delta P^{\mu}=\eta P\boldsymbol{y}^{\mu}\boldsymbol{y}^{\mu T}(I-P^{T}P)-\lambda P,$ (8)
where $I$ is the identity matrix. This rule adjusts $P$ incrementally so that its columns converge to the top $r$ principal components of the outputs $\{\boldsymbol{y}^{\mu}\}$.
By training both $Q$ and $P$, we allow the feedback matrix $B=QP$ to adaptively align with the relevant error directions, enabling the network to learn the correct mappings even under constrained feedback dimensionality.
Repeating the linear analysis that led to (5), we extend the derivation to our case with adaptive feedback weights. We define the transformed feedback matrix as $P=\bar{P}U^{T}$. This transformation aligns the feedback matrix $P$ with the principal components of the data, simplifying the analysis.
Taking the continuous-time limit (with $\eta\to 0$ and $\eta p=\tau$), we obtain a set of differential equations that describe the learning dynamics of the forward and backward weights:
$\tau\frac{d\bar{W}_{1}}{dt}$ $=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),$ (9)
$\tau\frac{d\bar{W}_{2}}{dt}$ $=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T}-\lambda\bar{W}_{2},$
and
$\tau\frac{dQ}{dt}$ $=\bar{W}_{1}(S-\bar{W}_{2}\bar{W}_{1})^{T}\bar{P}^{T}-\lambda Q,$ (10)
$\tau\frac{d\bar{P}}{dt}$ $=\bar{P}SS^{T}(I-\bar{P}^{T}\bar{P})-\lambda\bar{P}.$
Here, $\bar{B}=Q\bar{P}$ and $\lambda$ is the regularization parameter. Note that $Q$ is not affected by rotation, as it does not come in contact with either the input or the output. The full derivation can be found in the Appendix.
Notably, the updates to the feedback weights are local and follow learning rules that were well-studied in theoretical neuroscience [18, 19, 20, 21]. Local plasticity makes our framework an attractive alternative for backpropagation in brain circuit models.
We refer to this learning framework as low-dimensional feedback alignment (LDFA). Figure 2 compares the learning dynamics of LDFA with those of BP and FA where only the $Q$ matrix is learned. The results demonstrate that LDFA effectively aligns the feedback weights, enabling the network to converge to the correct solution despite the constrained error dimensionality. Furthermore, LDFA will learn the top $r$ components of $\Sigma$, even if $r<d$ (Figure 2e).
3.3 Restricted adaptive feedback in deep architectures
Our weight-update equations can be naturally extended to deeper networks. In the single hidden layer model above, the backward weights $P$ were updated using the true labels $\boldsymbol{y}^{\mu}$. However, in deeper models, hidden layers do not have ground-truth representations. Instead, each layer relies on the local error signal, which propagates through the network according to $\boldsymbol{\delta}_{l}^{\mu}=Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}$. This error signal $\boldsymbol{\delta}_{l}^{\mu}$ provides the necessary information for learning at the layer $l$.
To update the feedback weights $P_{l}$ in the absence of ground-truth representations, we use local error signals $\boldsymbol{\delta}_{l+1}^{\mu}$. As in the single-layer model, we use Oja’s rule to adjust $P_{l}$ to span the principal components of the error at the next layer, $\boldsymbol{\delta}_{l+1}^{\mu}$, ensuring efficient error propagation.
The complete update rules for layer $l$ are given by
$\Delta W_{l}^{\mu}&=\eta Q_{l}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{h}_{l}^{\mu T}-\lambda W_{l}\end{aligned}\qquad\text{and}\qquad\begin{aligned} \Delta Q_{l}^{\mu}&=\eta\boldsymbol{h}_{l}^{\mu}P_{l}\boldsymbol{\delta}_{l+1}^{\mu}-\lambda Q_{l},\\ \Delta P_{l}^{\mu}&=\eta P_{l}\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{\delta}_{l+1}^{\mu T}(I-P_{l}^{T}P_{l})-\lambda P,\end{aligned}$ (11)
where $\eta$ is the learning rate, $\lambda$ is the regularization parameter, $\boldsymbol{h}_{l}^{\mu}$ is the activation of layer $l$, and $\boldsymbol{\delta}_{l+1}^{\mu}$ is the error signal from the next layer.
By updating $P_{l}$ using the error signals, we ensure that the feedback weights of each layer are adapted to capture the most relevant directions in the error space, facilitating effective learning throughout the network. Notably, while we use the same learning rate $\eta$ and weight decay $\lambda$ for all components $\{W_{l}\}$, $\{Q_{l}\}$, and $\{P_{l}\}$, it can potentially differ.
In the interest of brevity, we omit simulations of deep linear networks, as the extension from the single-layer case is straightforward. Instead, we proceed directly to deep nonlinear networks, where the impact of constrained error feedback presents more complex and interesting dynamics.
4.   Nonlinear networks and complex architectures
Adapting our low-dimensional feedback alignment (LDFA) framework to nonlinear networks is straightforward because the core principles of local learning and constrained error feedback remain applicable. The local update rules in eq. (11) remain the same; the primary difference lies in the introduction of nonlinear activation functions during the propagation of signals and errors. Specifically, the forward and backward passes are modified as follows:
$h_{l}=f(W_{l}h_{l-1}),\quad\delta_{l}=Q_{l}P_{l}\delta_{l+1}\odot f^{\prime}(h_{l}),$ (12)
where $f$ is the nonlinear activation function applied element-wise, and $f^{\prime}$ is its derivative.
To test whether our learning rule extends effectively from linear to nonlinear models, we trained deep networks on the CIFAR-10 dataset. We used a simple nonlinear model with four fully connected layers of 512 ReLU neurons each. While not state-of-the-art, this model provides a suitable testbed for evaluating our theory’s applicability to nonlinear architectures and complex data.
To isolate the impact of feedback dimensionality, we applied the LDFA rule in eq. (11), constraining the rank of feedback matrices in one specific layer at a time while leaving the others unrestricted. We then measured the network’s test accuracy and compared it to backpropagation (BP) as a baseline. Figure 3a shows the accuracy as a function of the feedback rank $r_{l}$, with each curve representing a different constrained layer.
Consistent with our findings for the linear model, constraining the feedback rank to $r=d=10$, in any layer, match BP performance (Figure 3a), where $d$ is the number of classes. Interestingly, the shallower layers performed well even under tighter rank constraints ($r_{l}<d$), suggesting that the deeper layers compensate for the limited feedback in the earlier layers by effectively adjusting their weights.
This compensatory effect is possible because no information is lost during the feedforward pass, unlike in a bottlenecked network. To demonstrate that the network still utilizes high-dimensional representations, we compared the performance of LDFA-trained networks with constrained feedback to narrower networks without rank restrictions (Figure 3b). The results confirm that LDFA-trained networks leverage their width to maintain high performance despite feedback constraints.
Moreover, as shown in Figure 3b, nonlinear networks trained with low-rank feedback in all layers using LDFA can still match BP performance, demonstrating LDFA’s robustness even with dimensional constraints across the entire network.
Task dimensionality determines the minimal rank
Our linear analysis suggests that the error signal dimensionality needed for effective learning is tied to the loss gradient dimensionality, which depends on the number of classes in the data. To test this, we trained networks on subsets of CIFAR-100 with 50, 75, and 100 classes, constraining the ranks of all feedback matrices to $r$ (Figure 3c).
The results show that network performance matches BP when the feedback rank equals the number of classes ($r=d$). It indicates that the minimal rank required for effective learning aligns with the task’s complexity, defined by the number of output classes.
Low-dimensional error signal from the top layer
Our theory shows how to propagate error signals from deeper layers to shallower ones using restricted adaptive feedback. However, error projections can also bypass intermediate layers entirely, leading to different variants of direct low-dimensional feedbackalignment (dLDFA), analogous to Direct Feedback Alignment [22]. For example, we can make direct connections from the output or the penultimate layer to earlier layers, training these connections using our algorithm (Figure 3d). As with LDFA, this direct projection method matches the performance of BP when the rank of the feedback matrices satisfies $r\geq d$. This model is particularly important because it is more flexible and has greater potential to explain learning in the brain, where error signals may arrive from different pathways.
Figure 3: Low-dimensional feedback alignment (LDFA) efficiently trains nonlinear networks on CIFAR-10. (a) Constraining feedback in any layer to $r=d=10$ matches full BP performance. Different curves ($B_{1}$, $B_{2}$, and $B_{3}$) correspond to restricting feedback at different layers. (b) Reducing feedback dimensionality has minimal impact on performance, whereas reducing network width significantly degrades accuracy, indicating that high-dimensional representations are still utilized. (c) Minimal error dimensionality is sufficient for learning, as shown by subsampling classes from CIFAR-100. All feedback matrices are constrained to the same rank $r$. (d) Direct LDFA: all layers receive a low-dimensional error either directly from the output layer (solid) or from the penultimate layer (dashed), both converging to BP-level performance. (e) Performance of a 4-block VGG-like convolutional network trained with LDFA, constraining the layers with 512 channels. Inset: Training the same network with all layers constrained to a fraction of their size; the smallest layer has 64 channels. (f) A transformer-based architecture trained to classify images, demonstrating that LDFA generalizes to more advanced network architectures. See Supplemental Information for details on all implementations.
Convolutional neural networks
Our previous results demonstrate that fully connected networks can learn effectively from minimal error signals, even with complex datasets, matching backpropagation (BP) performance. Here, we extend this investigation to convolutional architectures.
Training convolutional networks with Feedback Alignment (FA) is notoriously difficult [8, 23]. Recent work has made progress by learning feedback weights within the FA framework [24], but our approach differs by restricting the error signal’s dimensionality using LDFA. We aim to determine whether convolutional networks can also benefit from low-dimensional error feedback.
We trained a VGG-like convolutional network [25] with four blocks and batch normalization on the CIFAR-10 dataset (see the appendix for details). Using LDFA, we decoupled error propagation from the feedforward pass in all layers. Initially, we constrained the feedback error only in the blocks containing 512 (Figure 3e). Consistent with our findings in fully connected networks, convolutional networks learn well with a feedback matrix with a rank similar to the number of classes, $d=10$.
To test whether convolutional networks can train when constraining all feedback paths, we further restricted each block to have feedback matrices with ranks equal to $1/2$, $1/4$, or $1/8$ of the block width (Figure 3e inset). Our results indicate that reducing the error dimensionality has a minimal impact on performance, except in the most extreme case. Specifically, constraining the feedback rank to $1/8$ of the block width resulted in a noticeable drop in performance. This finding aligns with our previous results, as the layer with 64 channels received feedback with a rank of $r=8$. Overall, our results demonstrate that convolutional networks can be efficiently trained using a minimal error signal.
Transformers
Transformers have become a dominant architecture in natural language processing and, more recently, have achieved state-of-the-art performance in image classification tasks [26, 27]. Unlike convolutional and standard feedforward architectures, transformers rely on a self-attention mechanism that computes interactions between tokens through the multiplication of key, query, and value matrices [28]. These operations introduce complex, nonlinear dependencies that make it unclear whether low-dimensional error feedback, as proposed in our framework, can effectively support learning in transformer models.

### v4

@@ 3.2 Training the feedback weights
As shown above, when $\mathrm{rank}(B)=r<m$ the stationarity condition in Eq. (6) is underdetermined, and fixed low-rank feedback (e.g., random $B$) can drive the dynamics to incorrect fixed points. We therefore keep the rank constraint, but allow the feedback pathway to adapt by parameterizing $B=QP$ with $Q\in\mathbb{R}^{k\times r}$ and $P\in\mathbb{R}^{r\times m}$.
The target feedback map is not known a priori, but backpropagation provides an ideal reference: exact credit assignment is obtained for $B=W_{2}^{\top}$. We thus learn the best rank-$r$ approximation of $W_{2}^{\top}$ by minimizing the factorization loss
$\mathcal{L}_{B}=\frac{1}{2}\|QP-W_{2}^{\top}\|_{F}^{2}.$ (8)
Notably, $\mathcal{L}_{B}$ involves only $W_{2}$ and the feedback parameters, so it is independent of the data and does not scale with batch size.
Gradient descent on $\mathcal{L}_{B}$ yields
$\Delta P$ $=Q^{\top}\bigl(W_{2}^{\top}-QP\bigr),$ (9)
$\Delta Q$ $=\bigl(W_{2}^{\top}-QP\bigr)P^{\top}.$
Thus, the feedback updates do not require access to activities or error signals; their only coupling to the training problem is through the evolving forward weights $W_{2}$.
As in the analysis above, we rotate into the singular basis of $\Sigma_{io}$ by writing $P=\bar{P}U^{\top}$ and $W_{2}=U\bar{W}_{2}$, where $U$ contains the left singular vectors of $\Sigma_{io}$. Taking the continuous-time limit (with learning rate $\eta_{B}=dt/\tau_{B}$) gives the mode-wise dynamics
$\tau_{B}\frac{d\bar{P}}{dt}$ $=Q^{\top}\bigl(\bar{W}_{2}^{\top}-Q\bar{P}\bigr),$ (10)
$\tau_{B}\frac{dQ}{dt}$ $=\bigl(\bar{W}_{2}^{\top}-Q\bar{P}\bigr)\bar{P}^{\top}.$
Together with the forward dynamics, Eqs. (10) define a closed system in which feedforward and feedback weights co-adapt. As shown in fig.˜2c,d, learning both $P$ and $Q$ eliminates the spurious low-rank fixed points and restores convergence to the correct solution under constrained feedback.
4.  Low-dimensional error in deep vision models
The linear theory above isolates the key obstruction induced by fixed low-rank feedback: when $\mathrm{rank}(B)=r<m$, the stationarity condition becomes underdetermined and learning can converge to spurious fixed points. Learning the feedback weights removes this degeneracy by making the backward mapping track the evolving feedforward transposes. Here we show that the same mechanism supports effective training in modern deep vision architectures, and we quantify the resulting scaling and computational consequences.
This extension is conceptually direct. The feedback factors are updated only through the mismatch $|Q_{l}P_{l}-W_{l}^{\top}|_{F}^{2}$, which depends on the feedforward weights but not on activities, batch statistics, or the sampled errors used to train $W_{l}$. Consequently, inserting local nonlinearities in the forward pass does not alter the feedback-learning rule. We first state the multilayer nonlinear construction explicitly, and then apply it to convolutional and transformer-based vision models.
4.1 Multilayer nonlinear perceptrons
We consider an $L$-layer multilayer perceptron with activities $\boldsymbol{h}0=\boldsymbol{x}$ and
$\boldsymbol{a}_{l+1}=W_{l}\boldsymbol{h}l,\hskip 18.49988pt\boldsymbol{h}_{l+1}=f(\boldsymbol{a}_{l+1}),\hskip 18.49988ptl=0,\ldots,L-1,$ (11)
where $f$ is applied element-wise. Let $\boldsymbol{\delta}_{l}$ denote the backpropagated error at layer $l$, including the derivative of the nonlinearity. Under low-dimensional feedback, we replace the transpose $W_{l}^{\top}$ in the backward recursion with a learned rank-$r$ factorization $B_{l}=Q_{l}P_{l}$, giving
$\boldsymbol{\delta}_{l}=(Q_{l}P_{l}\boldsymbol{\delta}_{l+1})\odot f^{\prime}(\boldsymbol{a}_{l}),\hskip 18.49988ptl=1,\ldots,L-1,$ (12)
with the usual initialization at the output layer determined by the loss.
Given these error signals, the feedforward weights follow the standard local outer-product update,
$\Delta W_{l}^{\mu}=\eta\boldsymbol{\delta}_{l+1}^{\mu}\boldsymbol{h}_{l}^{\mu\top}-\lambda W_{l},$ (13)
where $\lambda$ is an $L_{2}$ regularization (weight decay) parameter.
Finally, the feedback factors are trained to track the evolving transposes by gradient descent on $\frac{1}{2}|Q_{l}P_{l}-W_{l}^{\top}|_{F}^{2}$, yielding
$\Delta P_{l}=$ $Q_{l}^{\top}\big(W_{l}^{\top}-Q_{l}P_{l}\big),\qquad\text{and}$ (14)
$\Delta Q_{l}=$ $\big(W_{l}^{\top}-Q_{l}P_{l}\big)P_{l}^{\top}.$
These updates depend only on $W_{l}$ and can therefore be computed without storing activations or using training data; in practice, they may also use a separate step size and can be applied on a slower update schedule than the forward weights.
4.2 Deep convolutional networks
Training convolutional networks with Feedback Alignment (FA) is known to be challenging [4, 35]. Recent work improves FA by learning feedback weights [3], but our approach leverages a low-rank parameterization $B=QP$ that adapts both the output subspace that carries the propagated error (through $Q$) and the input subspace from which that error is assembled (through $P$). This factorized learning rule makes it possible to test whether deep convolutional networks can be trained with an error signal whose dimension is far smaller than the feedforward width, here measured by the number of channels per layer.
We trained a VGG-like convolutional network [63] with four blocks and batch normalization on CIFAR-10 (see Appendix for details). In all layers, we replaced transpose feedback with learned low-rank backward matrices. Within each block, we fixed the feedback rank to be a fraction of the block width: $1/2$, $1/4$, or $1/8$ (fig.˜3a). Reducing the error dimensionality had little effect on test performance, except in the most extreme setting. When the rank was $1/8$ of the block width, performance dropped noticeably, consistent with the linear prediction that accuracy degrades when $r<d$: in our smallest block (64 channels), rank-$1/8$ feedback corresponds to $r=8$, below the CIFAR-10 task dimensionality $d=10$. Overall, these results show that deep convolutional networks can be trained effectively using a strongly compressed error signal.
4.3 Transformer-based architectures
Transformers achieve state-of-the-art performance in image classification [18, 32]. Unlike convolutional and standard feedforward networks, they rely on self-attention, which mixes token representations through multiplicative interactions between key, query, and value projections [68]. These interactions introduce strong cross-token dependencies, so it is not a priori clear that strongly compressed error signals provide sufficient credit assignment for learning in transformer models.